OpenStudy (anonymous):

Find the minimum product of two numbers whose difference is 12. x-y=12

OpenStudy (anonymous):

Let x and y be the numbers. x-y=12 Product = xy = x(12-x) =12x-x^2 P=12x-x^2 dP/dx = 12-2x=0 2x=12 x=6 6-y=12 y=-6 minimum product = -36 Difference = 6-(-6)=12 The numbers are 6 and -6

OpenStudy (anonymous):

thanks alot!